# MagicFace: High-Fidelity Facial Expression Editing with Action-Unit Control

### v2

@@ III-C AU dropout
To improve the controllability of our method, we incorporate an AU dropout operation during training, enabling classifier-free guidance [32] for AU conditions. This strategy has been widely applied in conditional image generation to balance the trade-off between quality and diversity in images generated by latent diffusion models. In our training process, we randomly drop the AU as all zeros $\mathbf{c}_{uncon\_AU}=\varnothing$. To distinguish the dropped AUs from vectors representing no AU changes, we add a small amount of Gaussian noise to the vectors representing no AU changes, denoted as $\mathbf{c}_{zero\_AU}^{\prime}=\mathbf{c}_{zero\_AU}+\epsilon,\epsilon\sim\mathcal{N}(\mu,\sigma^{2})$. We empirically set $\sigma=0.2$. At the inference stage, we use a guidance scale $\alpha>1$ to modify the intensity of conditional control on the predicted noise:
@@ III-D Training Strategy
As introduced in Section III-B, the background and pose are formed into an independent image condition, so in the training stage we use the background of the target image parsed by [33]. We draw the pose contour on the background image by highlighting the 14 landmarks [34] around the chin of the face. Then in the inference time for edit, we use the background and pose parsed from the identity image to enable the generated results consistent with the identity in terms of background. We initialize the denoising UNet and the ID encoder with the pre-trained weights from SD. Both the weights of the denoising UNet and ID encoder are updated independently during training. Note that the ID encoder is only an encoder without time conditions, which needs only one step forward process before the multiple denoising steps in inference.
Figure 4: Qualitative comparison for facial expression editing. Other methods exhibit shortcomings in preserving the identity, pose or background, and they are unable to support continuous control over the intensity of generated expressions, whereas our method excels in maintaining exceptional detail features and meanwhile allows flexible, fine-grained control to the expression intensity. Please zoom in for more detailed observation.
@@ IV-A Implementations
Experimental settings. We collect 30K image pairs from the Aff-Wild dataset [35] by decomposing the videos into frames to train our model. Each pair has the same identity but different expressions, wherein at least one of the pose or background are highly likely different by ensuring that two sampled pairs maintain a certain distance in the video. We use LibreFace [9] to compute 12 AUs variation of the pair images and use Stable Diffusion 1-5 base  [^1] https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 to initialize the weights. Experiments are conducted on 4 NVIDIA A100 GPUs. Training is conducted for 100,000 steps with a batch size of 2 on each GPU. The learning rate is set to 1e-5.
Test details. For the test set, we edit 20 identities with variations on individual AUs and on AU combinations. We use 5 intensity variation levels for each single AU out of 12 AUs, and for AU combinations we use in total 387 combinations. This leads to a total number of $20\times(12\times 5+387)=8940$ edited images.
@@
Evaluation Criteria. We assess the performance of MagicFace by examining the edited image from four perspectives: the accuracy of AU intensity, identity preservation, background preservation and head pose preservation. We measure the AU values of the edited image $Y_{est}$ and compare it with the intended target values $Y_{tar}$ by computing the mean squared error (MSE). LibreFace [9] is used to estimate AU intensities of the generated images. To assess the similarity of identity, we measure the distance of embeddings between the edited image $\boldsymbol{I}_{tar}$ and the identity image $\boldsymbol{I}_{id}$, where the embedding is extracted from a pre-trained ID recongnition model  [^2] https://github.com/ageitgey/face_recognition. For background preservation, we employ the root pixel-wise mean squared error (RMSE) between the edited image and the identity image, where the face area of the both is masked to eliminate the interference of face expression difference. For head pose preservation, we compute head pose coefficients of the FLAME model [36] and compute their root mean squared error (RMSE).
@@
TABLE I: Comparison of methods in terms of AU accuracy, identity preservation and image similarity. N/A denotes incomputable. The best and second best results are reported in bold and [square brackets], respectively.
Method AU ID back- ground head pose
DeltaEdit [37] N/A 0.743 0.126 0.075
@@
DiffAE [15] N/A 0.579 0.161 0.183
GANmut [38] N/A [0.495] 0.059 0.032
FaceAdapter [39] [0.427] 0.567 [0.056] [0.034]
@@
IV-B Results
We compare our method against SoTA methods related with facial expression editing, including DeltaEdit [37], FaceEdit [^3] https://github.com/ototadana/sd-face-editor?tab=readme-ov-file, DALLE-3444https://openai.com/index/dall-e-3/, DiffAE [15], GANmut [38] and FaceAdapter [39]. DeltaEdit and FaceEdit cannot understand AU descriptions but they understand emotional vocabulary, so when testing on them, we use specific emotional words like happiness and adjectives that describe the intensity of expressions such as mild happiness, intense happiness in the text prompt. DALLE-3 is better at understanding AU text prompts and we design the prompt as add intensity level 2 (level range is [1, 10]) of AU1 , AU4 and AU15 to this man’s face which enables it to generate results as coherent to AU intensity as possible. Two emotional labels, i.e., happiness and sadness are employed respectively as the key words in text prompts. DiffAE provides the manipulation levels of smiling, so we use its positive direction as happiness and negative direction as sadness. GANmut is a facial expression model based on Valence-Arousal conditions. We employ the mapping between emotions and VA reported in their work to produce targeted expressions. FaceAdapter is designed to achieve face swapping and reenactment, we modify the condition of their model to enable only expression variation of the generated images. For fair comparison with these methods, we use the activated AUs related with these two labels [17] to test our model, which is AU6+AU12 for happiness and AU1+AU4+AU15 for sadness. The intensity of AUs grows gradually from 1 to 9 gapped by 2 from left to right in each row of each emotional label. Quantitative and qualitative results are shown in Tab. I and Fig. 4, respectively.
Comparisons. Our model surpasses the majority of existing methods in overall generation quality. Some approaches like DeltaEdit, DELLA-3 and FaceAdapter are unable to effectively preserve the identity of the input identity image. DELLA-3 performs well in continuously increasing the expression intensity and it can also interpret AU semantic meanings, but it can only generate the image with similar style to the input image, where the identity, background and pose are likely to be changed. Some methods including DiffAE, GANmut, DeltaFace are unable to continuously control the intensity of expression especifically for the sadness. DiffAE requires to crop and align the input image, which fails to preserve head pose. The results generated by FaceEdit are the most similar to ours, but they fail to preserve facial features well enough. Moreover, FaceEdit is unable to support the editing of other customed facial expressions.
Tab. I provides quantitative demonstration of the comparison, where we can observe that DELLA-3 and our model obtain much lower error in terms of the AU intensity accuracy. FaceEdit and GANmut show better score in terms of identity preservation as well as image similarity, corresponding to the observation from Fig. 4. Overall, our model significantly outperforms current either GAN-based or diffusion-based models with more precise and regulatable editing to facial expressions, both qualitatively and quantitatively.
@@ IV-B Results
Figure 5: Results of using laboratory dataset for training. In lab setting, the model cannot generalize to image from natural settings (the second row).
@@ IV-C Ablation Study
Training data. We investigate the effectiveness of using in-the-wild data to train our model by comparing with the model trained on DISFA [40]. DISFA provides AU annotation manually annotated by trained people, so literally using this dataset to train could produce more accurate AU condition than using AU conditions from AU estimation tools. However, as shown in Fig. 5, training on lab data can only generalize to new data in laboratory settings and cannot generalize to data from natural settings. The reason of this can be arrtibuted to the overly simplistic data patterns: simple background, only frontal face, and limited number of individuals in DISFA. Although AU estimator may bring inaccuray in estimated AU intensity of the training data, we can observe in Fig. 1 and 4 that it stills can generate edited images with high coherence to the AU prompt.
TABLE II: Ablation study on model architecture. The best and second best results are reported in bold and [square brackets], respectively.
AU ID back- ground head pose
@@
TABLE III: Ablation study on AU encoders. The best and second best results are reported in bold and [square brackets], respectively.
AU ID back- ground head pose
@@
Model design. To demonstrate the effectiveness of our model’s design, we explore four alternatives, 1) Replacing the whole architecture with ControlNet [14] (Control.) and add our Attribute Controller to enable background embedding; 2) Copy the architecture of 1) but enable denoising trainable (Cont. All); 3) Replacing the ID Encoder by a Conv layer (Conv ID.); 4) Replacing the ID Encoder by a CLIP image encoder [27] (CLIP ID.). Results are shown in Fig. 6 and Tab. II. Our design achieves the optimal performance. Design 1) and 2) shows that ControlNet can not adapt to AU conditions, and it significantly alters the background and the appearance of the face. Design 2) and 4) show that using CLIP or Conv to extract features can preserve image similarity but fails to preserve facial details. This further prevents them from precisely editing facial expressions based on AUs as some face details are lost. The edited results of these two experiments also differ from the identity image in some color details, like the hoodie and face color of the leftmost identity. We can also observe that in all of these ablations head pose is well-preserved, which indicates the effectiveness of separating background and pose as a pixel-aligned condition to make the model focus on the face area.
@@
Out-of-domain editing. It is worth noting that MagicFace demonstrates remarkable generalization to out-of-domain identity images with unseen AUs and characters, achieving impressive appearance controllability even without additional fine-tuning on the target domain. In the training stage the utmost AU variation given a neutral expression is ±$\pm 5$ (AU annotation only in range [0, 5]), but our model can edit the range of AUs almost extending from -10 to 10. This indicates that MagicFace has a strong capability to generate and explore contradictory and extreme expressions, as we can set AUs associated with contradictory emotional states to have opposite intensity variations and large values. On the other hand, on zero-shots results of applying our model to cartoon characters whose visual style is distinct from the training data of the real-human, our model can still ensure AU coherence with high-quality generation. We visualize the qualitative results in Fig. 7.
@@
Figure 7: Demonstration of out-of-domain testing for extreme expressions as well as unseen styles. Some AUs are out of the range [0, 5]. The left column displays the editing results of real person photos with extreme expressions and the right column displays that of the cartoon characters.
@@
AU Encoder. We investigate the influence of AU encoder by conducting two additional experiments: 1) Apply a MLP to encode AU vectors and concatenate it to the first Conv layer of the denoising UNet (MLP+Conv). 2) Append zero values to the end of the AU prompt to make the input vector of the same size as the time embedding and add it to the time embedding (ZeroAppend+Time). Quantitative results are shown in Tab. III. As can be observed, the encoding method of AUs does not significantly affect the performance of the approach, so we choose an easier way by just applying a linear layer to the AU prompt and then add it to the time embedding (Linear+Time).
AU dropout. We investigate the effects of AU dropout as well as the different values of the guidance scale $\alpha$. We first train our model without/with AU dropout. Then, for the model trained with AU dropout, we set $\alpha=0.5,1.5,2.5,3.0,3.5,4.5,5.5,6.5,7.5,8.5,9.5,10.5,11.5$, $12.5,13.5$ for classifier-free guidance. During inference, we ensure that all other parameters including the random seed remain consistent to allow for a fair comparison. As illustrated in Fig. 8, without AU dropout, classifier-free guidance is not supported, resulting in the worst generation quality where the target AU4[-6] is not obvious. For the model trained with dropout, when $\alpha>1.0$, the opposite variation of AU4 related area (brow lower towards brow higher) become progressively more distinct as $\alpha$ increases. However, color distortion will appear when $\alpha\geq 4.5$ and becomes increasingly noticeable starting from $\alpha=5.0$, as seen in the hoodie color, more evident beard and more red face. Tab. IV provides quantitative evidence supporting the effectiveness of our AU dropout. It also indicates that the optimal guidance scale is typically in the range of 1.5 to 3.5 in most cases. According to this study, we empirically set $\alpha=3.0$ for all test experiments.
@@
Figure 9: Visualization of self-attention maps (lst row) and cross-attention maps (2nd row) from the denoising UNet. Please zoom in to observe more details.
TABLE IV: Ablation study of AU dropout and different guidance scale values. The best and second best results are reported in bold and [square brackets], respectively. ”BG” is the abbreviation for ”background” and ”HP” is the abbreviation for ”head pose.” The best results are reported in bold.
@@
Visualization. Fig. 9 visualizes the attention maps learned in the self-attention as well as cross-attention of the denoising UNet. For the self-attention, we observe that initially the attention weights are almost uniformly distributed across the whole image, and then it focus more on the background area due to the input condition from Attribute Controller. As the denoising process nears its end, the attention gradually expands to include the facial regions. For the cross-attention, the attention shifts from an initially near-uniform distribution to gradually focusing on the facial regions, especially on the areas related with inputted AU varations. This indicates the self-attention layers mainly focus on the features of the image itself, while the cross attention focus more on interpreting AUs with the edited image.

### v3

@@ III-C AU dropout
To improve the controllability of our method, we incorporate an AU dropout operation during training, enabling classifier-free guidance [ho2021classifier] for AU conditions. This strategy has been widely applied in conditional image generation to balance the trade-off between quality and diversity in images generated by latent diffusion models. In our training process, we randomly drop the AU as all zeros $\mathbf{c}_{uncon\_AU}=\varnothing$. To distinguish the dropped AUs from vectors representing no AU changes, we add a small amount of Gaussian noise to the vectors representing no AU changes, denoted as $\mathbf{c}_{zero\_AU}^{\prime}=\mathbf{c}_{zero\_AU}+\epsilon,\epsilon\sim\mathcal{N}(\mu,\sigma^{2})$. We empirically set $\sigma=0.2$. At the inference stage, we use a guidance scale $\alpha>1$ to modify the intensity of conditional control on the predicted noise:
@@ III-D Training Strategy
As introduced in Section III-B, the background and pose are formed into an independent image condition, so in the training stage, we use the background of the target image parsed by [yu2021bisenet]. We draw the pose contour on the background image by highlighting the 14 landmarks [bulat2017far] around the chin of the face. Then, in the inference time for editing, we use the background and pose parsed from the identity image to enable the generated results to be consistent with the identity in terms of background. We initialize the denoising UNet and the ID encoder with the pre-trained weights from SD. Both the weights of the denoising UNet and ID encoder are updated independently during training. Note that the ID encoder is only an encoder without time conditions, which needs only one step forward process before the multiple denoising steps in inference.
@@ IV-A Implementations
Experimental settings. We collect 30K image pairs from the Aff-Wild dataset [kollias2019deep] by decomposing the videos into frames to train our model. Each pair has the same identity but different expressions, wherein at least one of the poses or backgrounds is highly likely to be different by ensuring that two sampled pairs maintain a certain distance in the video. We use LibreFace [chang2024libreface] to compute 12 AUs variation of the pair images and use Stable Diffusion 1-5 base  [^1] https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 to initialize the weights. Experiments are conducted on 4 NVIDIA A100 GPUs. Training is conducted for 100,000 steps with a batch size of 2 on each GPU. The learning rate is set to 1e-5.
@@
Evaluation Criteria. We assess the performance of MagicFace by examining the edited image from four perspectives: the accuracy of AU intensity, identity preservation, background preservation, and head pose preservation. We measure the AU values of the edited image $Y_{est}$ and compare them with the intended target values $Y_{tar}$ by computing the mean squared error (MSE). LibreFace [chang2024libreface] is used to estimate AU intensities of the generated images. To assess the similarity of identity, we measure the distance of embeddings between the edited image $\boldsymbol{I}_{tar}$ and the identity image $\boldsymbol{I}_{id}$, where the embedding is extracted from a pre-trained ID recognition model  [^2] https://github.com/ageitgey/face_recognition. For background preservation, we employ the root pixel-wise mean squared error (RMSE) between the edited image and the identity image, where the face area of both is masked to eliminate the interference of face expression difference. For head pose preservation, we compute head pose coefficients of the FLAME model [li2017learning] and compute their root mean squared error (RMSE).
Figure 4: Qualitative comparison for continuous facial expression editing. Our method excels in maintaining exceptional detail features of the face, while allowing flexible, fine-grained control over the expression intensity. Please zoom in for a more detailed observation.
IV-B Results
Continuous Facial Expression Editing. We edit 20 identities with variations on individual AUs and on AU combinations. We use 5 intensity variation levels for each single AU out of 12 AUs, and for AU combinations, we use an in total of 387 combinations. This leads to a total number of $20\times(12\times 5+387)=8940$ edited images. Then we compare our method with several representative approaches related to facial expression editing, including DeltaEdit [lyu2023deltaedit], FaceEdit [^3] https://github.com/ototadana/sd-face-editor?tab=readme-ov-file, DALLE-3444https://openai.com/index/dall-e-3/, DiffAE [preechakul2021diffusion], and GANmut [d2021ganmut]. For qualitative comparison, we show results on two emotions, i.e., happiness and sadness. For DeltaEdit and FaceEdit, which do not support AU descriptions, we use text prompts that contain emotional keywords together with adjectives describing the strength of the expression, for example, mild happiness and intense happiness. DALLE-3 can understand AU instructions, so we design prompts such as add intensity level 2 (level range [1, 10]) of AU1, AU4, and AU15 to this face in order to produce results that align with AU intensity as much as possible. DiffAE provides a one-dimensional control for smiling. Its positive direction is used to represent happiness, and its negative direction is used to represent sadness. GANmut is conditioned on Valence and Arousal. We follow the emotion-to-VA mapping provided in their work to generate the target expressions. In our method, we adopt the AUs associated with the two emotions following [ekman1978facial]. Specifically, AU6+AU12 is for happiness, and AU1+AU4+AU15 is for sadness. Fig. 4 presents the comparisons for the happiness and sadness expressions, and TABLE I summarizes the quantitative results for all editing settings.
Our model surpasses the majority of existing methods in overall generation quality. Approaches like DeltaEdit and DELLA-3 are unable to effectively preserve the identity of the input image. DELLA-3 performs well in continuously increasing the expression intensity, and it can also interpret AU semantic meanings, but it can only generate the image with a similar style to the input image, where the identity, background, and pose are likely to be changed. Some methods, including DiffAE, GANmut, DeltaFace, are unable to continuously control the intensity of expression specifically for sadness. The results generated by FaceEdit are the most similar to ours, but they fail to preserve facial features well enough. Moreover, FaceEdit is unable to support the editing of other customized facial expressions. TABLE I provides a quantitative demonstration of the comparison, where we can observe that DELLA-3 and our model obtain much lower error in terms of the AU intensity accuracy. FaceEdit and GANmut show better scores in terms of identity preservation as well as image similarity, corresponding to the observation from Fig. 4. Overall, our method provides strong capability and flexibility in controlling the intensity and continuity of facial expressions.
@@ IV-B Results
TABLE I: Comparison of methods in terms of AU accuracy, identity preservation, and image similarity. N/A denotes incomputable. The best and second best results are reported in bold and [square brackets], respectively.
Method AU ID
back-
ground
head
pose
MSE ($\downarrow$) L2 ($\downarrow$) RMSE ($\downarrow$) RMSEs ($\downarrow$)
DeltaEdit [lyu2023deltaedit] N/A 0.743 0.126 0.075
@@
DiffAE [preechakul2021diffusion] N/A 0.579 0.161 0.183
GANmut [d2021ganmut] N/A [0.495] 0.059 0.032
@@
TABLE II: Quantitative comparison of discrete expression editing performance. The best and second best results are reported in bold and [square brackets], respectively.
Method Emotion ID background head pose
Acc (%) $\uparrow$ L2 ↓ RMSE ↓ RMSE ↓
GANmut [lyu2023deltaedit] [90] 0.497 0.057 0.035
EmoStyle [azari2024emostyle] 63 0.549 0.071 0.087
InstantDrag [shin2024instantdrag] 66 [0.474] 0.051 0.069
DragDiffusion [shi2024dragdiffusion] 27 0.480 0.048 0.038
SDE-Drag [d2021ganmut] 46 0.495 [0.050] [0.037]
Step1X-Edit [liu2025step1x] 80 0.613 0.048 0.038
MagicFace (Ours) 92 0.466 0.048 [0.037]
Figure 5: Qualitative comparison with representative methods on discrete facial expression editing. The leftmost column shows the input images used as the editing source, and the remaining columns display the edited results for each target expression.
Figure 6: Qualitative comparison of facial expression transfer. The first column shows the source identities, and the second column displays the target expressions. AniPortrait, FaceAdapter, X-Portrait, and X-NeMo are face-reenactment methods, so their outputs inherit both the target expression and its pose.
Discrete Facial Expression Editing. We further evaluate our method on discrete facial expression editing. To make the comparison more comprehensive, we extend the set of discrete expressions beyond happiness and sadness and include three additional emotions, namely fear, surprise and disgust. We compare our model with several representative methods across different editing paradigms: activation–arousal based methods (GANmut [d2021ganmut], EmoStyle [azari2024emostyle]), dragging-based editing methods (DragDiffusion [shi2024dragdiffusion], InstantDrag [shin2024instantdrag], SDE-Drag [nie2023blessing]), and a recent text-driven method Step1X-Edit [liu2025step1x]. We randomly sample 20 test images and evaluate the editing performance from four perspectives: emotion accuracy, identity similarity, background consistency and head pose consistency. Emotion accuracy is estimated using the pretrained EmoNet classifier [toisoul2021estimation], while the other three metrics follow the evaluation procedures described in Sec. IV-A.
The qualitative and quantitative results are reported in Fig. 5 and TABLE II. As shown, our method consistently achieves the highest emotion accuracy across all discrete expressions, demonstrating its more precise control over the intended emotional changes. Furthermore, our approach preserves facial identity noticeably better than competing methods. Text-driven approaches such as Step1X-Edit [liu2025step1x] tend to unintentionally alter identity, while dragging-based methods may introduce pose inconsistencies. InstantDrag is particularly prone to such artifacts as it does not apply any masking mechanism. For example, when editing expressions such as happy, it may interpret the dragging direction as a head movement, leading to undesirable changes in pose (see the third row in Fig. 5). In addition, our method produces more realistic and fine-grained expression details. As illustrated in Fig. 5, especially in the fear and surprise cases, our results exhibit subtle cues such as forehead wrinkles that commonly appear in natural expressions. These details are not observed in GANmut [lyu2023deltaedit], even though it can generate expressions with relatively strong intensity.
TABLE III: Quantitative comparison of facial expression transfer performance. The best and second best results are reported in bold and [square brackets], respectively.
Method Expression ID AU
EXP-SIM $\uparrow$ L2 ↓ MSE ↓
ControlFace [jang2025controlface] 0.47 0.488 0.482
DiffusionRig [ding2023diffusionrig] 0.45 0.485 0.502
AniPortrait [wei2024aniportrait] 0.41 [0.477] 0.535
FaceAdapter [han2024face] 0.52 0.565 0.431
X-Portrait [xie2024x] 0.54 0.524 0.392
X-NeMo [zhao2025x] 0.66 0.513 0.253
MagicFace (Ours) [0.61] 0.471 [0.288]
Facial Expression Transfer. Our method is further evaluated on a facial expression transfer task. Since facial expressions can be manipulated through various types of conditioning, several representative approaches across different paradigms are included for comparison. For 3D-based control, ControlFace [jang2025controlface] and DiffusionRig [ding2023diffusionrig] are selected. For landmark-based control, AniPortrait [wei2024aniportrait] and FaceAdapter [han2024face] are considered. For drive-image-based control, X-Portrait [xie2024x] and X-NeMo [zhao2025x] are evaluated. AniPortrait, FaceAdapter, X-Portrait, and X-NeMo are face reenactment methods in which expression and head pose are inherently entangled. Therefore, pose and background consistency are not evaluated in this setting. Instead, the comparison focuses on three aspects: the precision of expression transfer, identity similarity, and AU accuracy. The accuracy of expression transfer is measured using the pretrained emotion encoder EmoNet [toisoul2021estimation]. Specifically, the mean concordance correlation coefficient and Pearson correlation coefficient of valence and arousal are computed, and an overall emotion similarity score (EXP-SIM) is obtained. This metric reflects how precisely fine-grained expression variations are captured.
The qualitative and quantitative comparisons are presented in Fig. 6 and TABLE III. Our method achieves the strongest identity preservation and obtains competitive expression transfer accuracy. Although X-NeMo reaches higher EXP-SIM values, visual inspection reveals several limitations. As shown in the fourth row of Fig. 6, X-NeMo may inherit identity traits from the driving image, resulting in noticeable identity leakage, and tends to distort the head size of the source subject (other rows). In contrast, our method consistently maintains the identity characteristics of the source while transferring the target expression. Moreover, compared with 3D-based and landmark-based methods, our approach better captures subtle and fine-grained expression cues. For instance, in rows three and four of Fig. 6, expressions involving frowning or widening of the eyes naturally produce forehead wrinkles. These details are faithfully reproduced by our method but are missing in the 3D-based and landmark-based approaches, which have limited representational capacity for such nuances. This highlights the advantage of AU-based conditioning for achieving both precise and expressive facial animation.
@@
Figure 7: Results of using the laboratory dataset for training. In a lab setting, the model cannot generalize to images from natural settings (the second row).
@@ IV-C Ablation Study
Training data. We investigate the effectiveness of using in-the-wild data to train our model by comparing with the model trained on DISFA [mavadati2013disfa]. DISFA provides AU annotation manually annotated by trained people, so literally using this dataset to train could produce more accurate AU conditions than using AU conditions from AU estimation tools. However, as shown in Fig. 7, training on lab data can only generalize to new data in laboratory settings and cannot generalize to data from natural settings. The reason for this can be attributed to the overly simplistic data patterns: simple background, only frontal face, and a limited number of individuals in DISFA. Although the AU estimator may bring inaccuracies in the estimated AU intensity of the training data, we can observe in Fig. 1 and 4 that it can still generate edited images with high coherence to the AU prompt.
TABLE IV: Ablation study on model architecture. The best and second best results are reported in bold and [square brackets], respectively.
AU ID
back-
ground
head
pose
@@
TABLE V: Ablation study on AU encoders. The best and second best results are reported in bold and [square brackets], respectively.
AU ID
back-
ground
head
pose
@@
Model design. To demonstrate the effectiveness of our model’s design, we explore four alternatives, 1) Replacing the whole architecture with ControlNet [zhang2023adding] (Control.) and add our Attribute Controller to enable background embedding; 2) Copy the architecture of 1) but enable denoising trainable (Cont. All); 3) Replacing the ID Encoder by a Conv layer (Conv ID.); 4) Replacing the ID Encoder by a CLIP image encoder [radford2021learning] (CLIP ID.). Results are shown in Fig. 8 and TABLE IV. Our design achieves the optimal performance. Design 1) and 2) show that ControlNet can not adapt to AU conditions, and it significantly alters the background and the appearance of the face. Design 2) and 4) show that using CLIP or Conv to extract features can preserve image similarity, but fail to preserve facial details. This further prevents them from precisely editing facial expressions based on AUs, as some facial details are lost. The edited results of these two experiments also differ from the identity image in some color details, like the hoodie and face color of the leftmost identity. We can also observe that in all of these ablations, head pose is well-preserved, which indicates the effectiveness of separating background and pose as a pixel-aligned condition to make the model focus on the face area.
@@
Out-of-domain editing. It is worth noting that MagicFace demonstrates remarkable generalization to out-of-domain identity images with unseen AUs and characters, achieving impressive appearance controllability even without additional fine-tuning on the target domain. In the training stage, the utmost AU variation given a neutral expression is $\pm 5$ (AU annotation only in range [0, 5]), but our model can edit the range of AUs, almost extending from -10 to 10. This indicates that MagicFace has a strong capability to generate and explore contradictory and extreme expressions, as we can set AUs associated with contradictory emotional states to have opposite intensity variations and large values. On the other hand, on zero-shot results of applying our model to cartoon characters whose visual style is distinct from the training data of the real-human, our model can still ensure AU coherence with high-quality generation. We visualize the qualitative results in Fig. 9.
@@
Figure 9: Demonstration of out-of-domain testing for extreme expressions as well as unseen styles. Some AUs are out of the range [0, 5]. The left column displays the editing results of real person photos with extreme expressions, and the right column displays those of the cartoon characters.
@@
AU Encoder. We investigate the influence of the AU encoder by conducting two additional experiments: 1) Apply an MLP to encode AU vectors and concatenate it to the first Conv layer of the denoising UNet (MLP+Conv). 2) Append zero values to the end of the AU prompt to make the input vector of the same size as the time embedding and add it to the time embedding (ZeroAppend+Time). Quantitative results are shown in TABLE V. As can be observed, the encoding method of AUs does not significantly affect the performance of the approach, so we choose an easier way by just applying a linear layer to the AU prompt and then adding it to the time embedding (Linear+Time).
@@
AU dropout. We investigate the effects of AU dropout as well as the different values of the guidance scale $\alpha$. We first train our model without/with AU dropout. Then, for the model trained with AU dropout, we set $\alpha=0.5,1.5,2.5,3.0,3.5,4.5,5.5,6.5,7.5,8.5,9.5,10.5,11.5$, $12.5,13.5$ for classifier-free guidance. During inference, we ensure that all other parameters, including the random seed, remain consistent to allow for a fair comparison. As illustrated in Fig. 10, without AU dropout, classifier-free guidance is not supported, resulting in the worst generation quality where the target AU4[-6] is not obvious. For the model trained with dropout, when $\alpha>1.0$, the opposite variation of AU4 related area (brow lower towards brow higher) becomes progressively more distinct as $\alpha$ increases. However, color distortion will appear when $\alpha\geq 4.5$ and becomes increasingly noticeable starting from $\alpha=5.0$, as seen in the hoodie color, more evident beard, and more red face. TABLE VI provides quantitative evidence supporting the effectiveness of our AU dropout. It also indicates that the optimal guidance scale is typically in the range of 1.5 to 3.5 in most cases. According to this study, we empirically set $\alpha=3.0$ for all test experiments.
TABLE VI: Ablation study of AU dropout and different guidance scale values. The best and second best results are reported in bold and [square brackets], respectively. ”BG” is the abbreviation for ”background”, and ”HP” is the abbreviation for ”head pose.” The best results are reported in bold.
@@
TABLE VII: Efficiency analysis of the training and inference stages of the proposed approach. Per iter. means per iteration.
Setting GPU Type dtype Memory Time (s)
Training A100 40GB ($\times 4$) FP32 26.5GB 23.6 (per iter.)
Inference A100 40GB FP32 8.2GB 8.4
Efficiency Analysis. The training and inference efficiency of our method is summarized in TABLE VII. The results show that the proposed framework exhibits reasonable computational cost for a large-model fine-tuning task, although it is inherently slower than non-diffusion-based approaches. Nevertheless, this additional cost brings the benefit of substantially finer control over facial expressions and higher-quality editing results. Training on four A100-40GB GPUs requires 26.5 GB of memory and approximately 23.6 seconds per iteration under FP32 precision, which is acceptable given the scale of the model. During inference, the model operates on a single A100-40GB GPU with a memory footprint of only 8.2 GB and a runtime of 8.4 seconds per image. These results demonstrate that the method is practical to deploy and provides efficient performance at test time.